| ID | 
Date | 
Author | 
Category | 
Subject | 
Text | 
  |  
| 
  250  
 | 
Sun Apr  6 22:53:53 2025 | 
Michael | Analysis | Added shift tracking and power threshold to Shahab's npz_addup.py | I added automatic shift tracking and a check 
for a minimum wanted power in that bin to 
Shahab's npz_addup.py script.
 | 
   |  | 
  249  
 | 
Sun Apr  6 22:22:46 2025 | 
Taka, Moriguchi | General | Lassie screenshots |   | 
       |  | 
  248  
 | 
Sun Apr  6 21:29:37 2025 | 
Taka, Moriguchi | Accelerator | Beam sharing | again, ESR : FRS = 1 : 1 | 
   |  | 
  247  
 | 
Sun Apr  6 21:12:07 2025 | 
Taka, Moriguchi | General | Lassie screenshots |   | 
       |  | 
  246  
 | 
Sun Apr  6 20:49:25 2025 | 
Ruijiu Chen | Analysis | Spectrum during April 6 20:00- 20:19 |   | 
       |  | 
  245  
 | 
Sun Apr  6 20:13:03 2025 | 
Taka, Moriguchi | Accelerator | Beam sharing | Now we have every injection. | 
   |  | 
  244  
 | 
Sun Apr  6 19:42:23 2025 | 
Taka, Moriguchi | General | Lessie screenshots |   | 
       |  | 
  243  
 | 
Sun Apr  6 19:23:09 2025 | 
Shahab, Moriguchi, Taka | Accelerator | Beam sharing | Beam sharing with FRS is now 1 to 1. | 
   |  | 
  242  
 | 
Sun Apr  6 18:48:36 2025 | 
Shahab, Moriguchi, Taka | Accelerator | Beam back | Beam is back now. | 
   |  | 
  241  
 | 
Sun Apr  6 18:42:51 2025 | 
Shahab, Taka, Moriguchi | General | Lassie screenshots |   | 
       |  | 
  240  
 | 
Sun Apr  6 18:36:56 2025 | 
Shahab, Moriguchi, Taka | Accelerator | Beam stopped | Beam stopped due to device trouble. | 
   |  | 
  239  
 | 
Sun Apr  6 17:29:04 2025 | 
Ruijiu Chen | Analysis | 188Hf72+ spectrum and single ion line from NTCAP data | This shows the spectrum of 188Hf72+ during 
"April06_01:00_02:00". Until April 6 2:00, 
the Bp = 7.918262 Tm has not been changed. | 
       |  | 
  238  
 | 
Sun Apr  6 17:24:47 2025 | 
Shahab, Taka, Moriguchi | General | Lassie screenshots |   | 
       |  | 
  237  
 | 
Sun Apr  6 16:24:36 2025 | 
Wolfram, Shahab, Moriguchi, Taka | DAQ | NTCAP restart | NTCAP restarted with new timestamp. | 
   |  | 
  236  
 | 
Sun Apr  6 16:16:22 2025 | 
Wolfram, Shahab, Taka, Moriguchi | General | Lassie screenshots |   | 
       |  | 
  235  
 | 
Sun Apr  6 15:13:47 2025 | 
Burcu, Moriguchi, Taka, Gregor | General | Lassie screenshots |   | 
       |  | 
  234  
 | 
Sun Apr  6 14:17:07 2025 | 
Ragan, Burcu, Gregor , Yuri, Taka | General | Every injection to the ESR | From 14:15, we have every injection to the 
ESR. FRS is on pause for 1-2 hours. | 
   |  | 
  233  
 | 
Sun Apr  6 13:54:25 2025 | 
Ragan, Burcu, Gregor | General | Lassie screenshots |   | 
       |  | 
  232  
 | 
Sun Apr  6 12:57:23 2025 | 
Ragan, Burcu, Gregor | General | Lassie screenshots |   | 
       |  | 
  231  
 | 
Sun Apr  6 11:54:45 2025 | 
Ragan, Burcu, Gregor | General | Lassie screenshots |   | 
       |   
 |